Find the product of the given complex number and its conjugate. Write the answer in the form a + b i.
1/3 -8i

Answer by Alan3354(69443)

Find the product of the given complex number and its conjugate. Write the answer in the form a + b i.
1/3 -8i
---------------
= (1/3 - 8i)*(1/3 + 8i)
= 1/9 - 64i^2
= 1/9 + 64
= 577/9 + 0i